# Congruences for the ratios of Rankin–Selberg $L$-functions

### v1

@@ Abstract.
A well-known principle states that a congruence between objects should give rise to a corresponding congruence between the special values of $L$-functions attached to these objects. In this article we computationally investigate this principle for Rankin–Selberg $L$-functions attached to pairs of holomorphic cuspforms.
@@ 1. Introduction
It is possible that the ratios of $L$-vaues are not integral in $F$; the congruence is to interpreted as the difference of the ratios has a positive $\mathfrak{P}$-adic valuation.
@@
The computations rely on two algorithms. The first algorithm, reviewed in 2.2, computes the special values of Rankin-Selberg $L$-functions, and is essentially due to Shimura and Hida. Roughly speaking, it relies on the interpretation of the $L$-value as a Petersson inner product of a cusp form with a holomorphic projection which is due to Shimura [15], and the computation of the holomorphic projection recursively from the given data and an appropriate Eisenstein series due to Hida [5]. The second well-known algorithm, reviewed in 2.3, computes the $L$-values of modular forms from the Hecke-equivariant pairing between a space of cusp forms and a suitable space of modular symbols. It should be mentioned that there is also an algorithm by Tim Dokchitser [2] to compute the $L$-values of modular forms numerically. All the computations in this paper are done on SAGE [14].
Based on these examples, we formulate a conjecture in Conj. 4.1. In the companion article [11], we prove a variation of this conjecture using the machinery of Eisenstein cohomology as developed by Harder and the second author in [3]. As the proofs in [3] involves a cohomological interpretation of Langlands’s theorem on the constant term of an Eisenstein series, the results and proofs in [11] are completely independent of the computational techniques this paper.
@@ 2.1. Review of some results of Shimura and Hida
Let $\Gamma_{1}(N)$ be the standard Hecke congruence subgroup of $\mathrm{SL}_{2}(\mathbb{Z})$; the volume of its fundamental domain $\Phi_{1}(N)$ with respect to the measure $y^{-2}dxdy$ will be denoted $\mu(\Phi_{1}(N)).$ Let $M_{k}(\Gamma_{1}(N))$ be the space of holomorphic modular forms of weight $k$ for $\Gamma_{1}(N)$. For $f,\,h\in M_{k}(\Gamma_{1}(N))$ such that $fh$ is a cusp form, define the Petersson inner product $\langle f,h\rangle$ by:
@@
where $L_{N}(s,\omega)$ denotes the Dirichlet $L$-series $\sum_{n=1}^{\infty}\omega(n)/n^{s}$ and the subscript $N$ means that with $\omega(n)=0$ if ${\rm gcd}(n,N)\neq 1.$ If $\omega$ is primitive modulo $N$ then this condition is automatic and the subscript may be dropped. One has the following integral representation:
@@
The Maaß–Shimura differential operators on the space of smooth functions on the Poincare upper half plane $\mathbb{H}$ by
@@
Let $A_{r}$ denote the set of all functions of the form $h(z)=\sum_{\nu=0}^{r}y^{-nu}g_{\nu}(z)$ with holomorphic functions $g_{\nu}$ on $\mathbb{H}$ which have Fourier expansions $g_{\nu}(z)=\sum_{n=0}^{\infty}b_{\nu n}q^{n}$. Elements of $A_{r}$ may be called nearly holomorphic modular forms. The following lemma is [15, Lem. 7].
@@
The following orthogonality relation is [15, Lem. 6].
@@
$\langle f_{\rho},\delta_{\lambda}^{(r)}E_{\lambda,N}\rangle\ =\ \langle f_{\rho},h_{0}\rangle.$
The following theorem is one of the main results from [15]:
@@
$D(k-1-r,f\times g)\ =\ c_{r}\pi^{k}\langle f_{\rho},\ g.\delta_{\lambda}^{(r)}E_{\lambda,N}^{*}(z,\chi\psi)\rangle,$
@@
The $h_{\nu}$, as in equation (3), satisfy the following recursive relations:
@@
    When $\lambda=2$ and $\chi\psi=\mbox{trivial}$ then
@@
    When $\lambda\neq 2$ or $\chi\psi\neq\mbox{trivial}$ then
@@
where $L_{\infty}(s,f\times g)=(2\pi)^{-2s}\Gamma(s)\Gamma(s+1-l)$ is the archimdean factor and for a Dirichlet character $\omega$ modulo $N$ put $L_{N}(s,\omega)=\sum_{n=1}^{\infty}\omega(n)n^{-s}$ with $\omega(n)=0$ if $(n,N)=1.$ If $g$ is also a normalized eigenform in $S_{l}(N,\psi),$ then $L(s,f\times g)$ satisfies a functional equation
@@
See Shimura [15, Section 3] and Hida [5, Section 9]. The line of symmetry for this functional equation is $\text{Re}(s)=(k+l-1)/2.$ An integer $m$ is critical for $L(s,f\times g),$ if the Gamma factors on both sides of the functional equation are finite at $s=m,$ i.e., if $\Gamma(m)\Gamma(m+1-l)$ and $\Gamma(k+l-1-m)\Gamma(k-m)$ are finite. Therefore the critical set is:
@@
Let $f$ be a normalized eigenform of $S_{k}(N)$, $g$ an element of $G_{l}(N)$, and $m$ a positive integer. Suppose $l<k$ and $\frac{1}{2}(k+l-2)<m<k$. Then $\pi^{-k}\langle f,f\rangle^{-1}D(m,f\times g)$ belongs to $\mathbb{Q}(f)\mathbb{Q}(g)$, the compositum of $\mathbb{Q}(f)$ and $\mathbb{Q}(g)$. Moreover, for every automorphism $\sigma$ of $\mathbb{C}$, we have
@@
The parameters determining the abelian part and the infinite part of $L(s,f\times g)$ in (4) depend only on the weights and the nebentype characters involved. So for two $f,f^{\prime}\in S_{k}(N,\chi)$ and $g\in M_{l}(N,\psi)$ one has
@@
(7) $\frac{L(m,\omega)}{L(m+2,\omega)}=(m+2)(m+1)\Bigl(\frac{N}{2\pi}\Bigr)^{2}\frac{B_{m,\bar{\omega}}}{B_{m+2,\bar{\omega}}}.$
If $\omega$ modulo $N$ is not primitive take $\omega^{\text{prim}}$ to be the primitive Dirichlet character defined by $\omega$. Then $L_{N}(s,\omega)=\prod_{p|N}\left(1-\omega^{\text{prim}}(p)p^{-s}\right)L(s,\omega^{\text{prim}}).$ Hence, from (7) and Theorem 2.7, for any integer $m$ with $\frac{k+l-2}{2}<m<m+1<k$, it follows that:
@@ 2.2. An algorithm for the algebraic part of $D(m,f\times g)$
    Extend $f_{\rho}$ to a basis consisting of eigenforms $S_{k}(\Gamma_{1}(N))$ which we denote by $\mathcal{B}=\{f_{0}:=f_{\rho},f_{1},\dots,f_{n}\}$. (The dimension is $n$.)
@@
This is the desired value due to Lemma 2.4, Lemma 2.3, Theorem 2.5, and $\langle f_{0},f_{i}\rangle=0$ for $i=1,\cdots,n$ and $\langle f^{\rho},f^{\rho}\rangle=\langle f,f\rangle$. The constant $c_{r}$ appears in Theorem 2.5.
@@
2.3. An algorithm for the algebraic part of $L(m,f)$
@@ 2.3.1. Special values of $L$-functions via modular symbols
$P,\gamma]:=\gamma(P\otimes\{0,\infty\})\in\mathbb{M}_{k}(\Gamma)$
When $\Gamma=\Gamma_{1}(N)$ and if $\gamma=\left(\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\right),\ \gamma^{\prime}=\left(\begin{smallmatrix}a^{\prime}&b^{\prime}\\ c^{\prime}&d^{\prime}\end{smallmatrix}\right)\in\Gamma_{1}(N)$ are such that $(c,d)\equiv(c^{\prime},d^{\prime})\pmod{N}$, then $P,\gamma]=[P,\gamma^{\prime}]$. Hence, the Manin symbol $P,\gamma]$ is determined by $P$ and the lower two entries $c,d$ of the matrix $\gamma$. So we take $[P,(c,d)]$ itself to be a Manin symbol. For the following theorem, see, for example, Stein [17, Prop. 8.3].
@@
The Manin Symbols generate $\mathbb{M}_{k}(\Gamma).$
One can define Hecke operators on the space of modular symbols; see Stein [17, Section 8.3].
@@
Let $S_{k}(\Gamma),\bar{S}_{k}(\Gamma)$ denote the space of holomorphic and anti-holomorphic cusp forms respectively. Then the pairing $(\cdot,\cdot):(S_{k}(\Gamma)\oplus\bar{S}_{k}(\Gamma))\times\mathbb{M}_{k}(\Gamma)\rightarrow\mathbb{C}$
is Hecke equivariant, i.e., for $x\in\mathbb{M}_{k}(\Gamma)$ we have $(T_{n}(f_{1},f_{2}),x)=(f_{1},f_{2},T_{n}(x)).$
For a cusp form $f\in S_{k}(N)$, the associated $L$-function is
(9) $D(s,f)=\sum_{n=1}^{\infty}\frac{a_{n}}{n^{s}},\quad{\rm and}\quad\int_{0}^{\infty}f(it)t^{s}\frac{dt}{t}=(2\pi)^{-s}\Gamma(s)D(s,f),\quad\Re(s)\gg 0.$
Consider the modular symbols $X^{j}Y^{k-2-j}\otimes\{0,\infty\}$ for $j=0,1,2,\dots,k-2$. Then the non-degenerate and Hecke equivariant bilinear pairing $(\cdot,\cdot)$ of these modular symbols against the cusp form $f$ is equal to
@@
Hence, by (9), we get the following relation between the bilinear pairing and the special value of $L$-function. For $m=1,2\dots,k-1$,
@@
For $m=1,2\dots,k-1$, define algebraic part of $L(m,f)$ by:
@@
2.3.2. The algorithm
Input: $f\in S_{k}(\Gamma_{1}(N))$ a newform. A prime $p$ for which $a(p,f)$ is also given.
Output: $\frac{D(m,f)}{(-2\pi\sqrt{-1})^{m-1}D(1,f)}$ for $m=1,3,\dots,$ and $\frac{D(m,f)}{(-2\pi\sqrt{-1})^{m-2}D(2,f)}$ for $m=2,4,\dots$.

### v2

@@ Abstract.
A well-known principle states that a congruence between objects should give rise to a corresponding congruence between the special values of $L$-functions attached to these objects. We computationally investigate this principle for Rankin–Selberg $L$-functions attached to pairs of holomorphic cuspforms, and formulate a precise conjecture in general.
Key words and phrases:
Congruences, Special Values, Rankin–Selberg $L$-functions
@@ 1. Introduction
It is possible that the ratios of $L$-vaues are not integral in $F$; the congruence is to be interpreted as the difference of the ratios has a positive $\mathfrak{P}$-adic valuation.
@@
The computations rely on two algorithms. The first algorithm, reviewed in 2.2, computes the special values of Rankin-Selberg $L$-functions, and is essentially due to Shimura and Hida. Roughly speaking, it relies on the interpretation due to Shimura [14] of the $L$-value as a Petersson inner product of a cusp form with a holomorphic projection, and the computation due to Hida [5] of the holomorphic projection recursively from the given data and an appropriate Eisenstein series. The second well-known algorithm, reviewed in 2.3, computes the $L$-values of modular forms from the Hecke-equivariant pairing between a space of cusp forms and a suitable space of modular symbols. It should be mentioned that there is also an algorithm by Tim Dokchitser [2] to compute the $L$-values of modular forms numerically. All the computations in this paper are done on SAGE [13].
Based on these examples, we formulate a conjecture in Conj. 4.1. In the companion article [10], we prove a variation of this conjecture using the machinery of Eisenstein cohomology as developed by Harder and the second author in [3]. As the proofs in [3] involve a cohomological interpretation of Langlands’s theorem on the constant term of an Eisenstein series, the results and proofs in [10] are completely independent of the computational techniques this paper.
@@ 2.1. Review of some results of Shimura and Hida
Let $\Gamma_{1}(N)$ be the standard Hecke congruence subgroup of $\mathrm{SL}_{2}(\mathbb{Z})$; the volume of its fundamental domain $\Phi_{1}(N)$ with respect to the measure $y^{-2}dxdy$ will be denoted $\mu(\Phi_{1}(N)).$ Let $M_{k}(\Gamma_{1}(N))$ be the space of holomorphic modular forms of weight $k$ for $\Gamma_{1}(N)$. Sometimes $M_{k}(\Gamma_{1}(N))$ is abbreviated as $M_{k}(N).$ For $f,\,h\in M_{k}(\Gamma_{1}(N))$ such that $fh$ is a cusp form, define the Petersson inner product $\langle f,h\rangle$ by:
@@
where $L_{N}(s,\omega)$ denotes the Dirichlet $L$-series $\sum_{n=1}^{\infty}\omega(n)/n^{s}$ and the subscript $N$ means that $\omega(n)=0$ if ${\rm gcd}(n,N)\neq 1.$ If $\omega$ is primitive modulo $N$ then this condition is automatic and the subscript may be dropped. One has the following integral representation:
@@
The Maaß–Shimura differential operators on the space of smooth functions on the Poincaré upper half plane $\mathbb{H}$ are defined by:
@@
Let $A_{r}$ denote the set of all functions of the form $h(z)=\sum_{\nu=0}^{r}y^{-\nu}g_{\nu}(z)$ with holomorphic functions $g_{\nu}$ on $\mathbb{H}$ which have Fourier expansions $g_{\nu}(z)=\sum_{n=0}^{\infty}b_{\nu n}q^{n}$. Elements of $A_{r}$ may be called nearly holomorphic modular forms. The following lemma is [14, Lem. 7].
@@
The following orthogonality relation is [14, Lem. 6].
@@
$\langle f_{\rho},g\cdot\delta_{\lambda}^{(r)}E_{\lambda,N}\rangle\ =\ \langle f_{\rho},h_{0}\rangle,$ where $h_{0}$ is the holomorphic projection of the nearly holomorphic modular form $g\cdot\delta_{\lambda}^{(r)}E_{\lambda,N}.$
Combining all of the observations from before we get
@@
$D(k-1-r,f\times g)\ =\ c_{r}\pi^{k}\langle f_{\rho},\ g.\delta_{\lambda}^{(r)}E_{\lambda,N}^{*}(z,\chi\psi)\rangle=\langle f_{\rho},h_{0}\rangle,$
@@
The functions $h_{\nu}$, as in (3), satisfy the following recursive relations:
@@
    When $\lambda=2$ and $\chi\psi=1\!\!1$ then
@@
    When $\lambda\neq 2$ or $\chi\psi\neq 1\!\!1$ then
@@
where $L_{\infty}(s,f\times g)=(2\pi)^{-2s}\Gamma(s)\Gamma(s+1-l)$ is the archimdean factor and for a Dirichlet character $\omega$ modulo $N$ put $L_{N}(s,\omega)=\sum_{n=1}^{\infty}\omega(n)n^{-s}$ with $\omega(n)=0$ if $(n,N)\neq 1.$ If $g$ is also a normalized eigenform in $S_{l}(N,\psi),$ then $L(s,f\times g)$ satisfies a functional equation
@@
See Shimura [14, Section 3] and Hida [5, Section 9]. The line of symmetry for this functional equation is $\text{Re}(s)=(k+l-1)/2.$ An integer $m$ is critical for $L(s,f\times g),$ if the Gamma factors on both sides of the functional equation are finite at $s=m,$ i.e., if $\Gamma(m)\Gamma(m+1-l)$ and $\Gamma(k+l-1-m)\Gamma(k-m)$ are finite. Therefore the critical set is:
@@
Let $f$ be a normalized eigenform of $S_{k}(N)$, $g$ an element of $G_{l}(N)$, and $m$ a positive integer. Suppose $l<k$ and $\lfloor\frac{1}{2}(k+l-1)\rfloor<m<k$. Then $\pi^{-k}\langle f,f\rangle^{-1}D(m,f\times g)$ belongs to $\mathbb{Q}(f)\mathbb{Q}(g)$, the compositum of $\mathbb{Q}(f)$ and $\mathbb{Q}(g)$. Moreover, for every automorphism $\sigma$ of $\mathbb{C}$, we have
@@
The parameters determining the abelian part and the infinite part of $L(s,f\times g)$ in (4) depend only on the weights and the nebentype characters involved. For $f,f^{\prime}\in S_{k}(N,\chi)$ and $g\in M_{l}(N,\psi)$ one has
@@
(7) $\frac{L(m,\omega)}{L(m+2,\omega)}\ =\ -(m+2)(m+1)\Bigl(\frac{N}{2\pi}\Bigr)^{2}\frac{B_{m,\bar{\omega}}}{B_{m+2,\bar{\omega}}}.$
If $\omega$ modulo $N$ is not primitive take $\omega^{\text{prim}}$ to be the primitive Dirichlet character defined by $\omega$. Then $L_{N}(s,\omega)=\prod_{p|N}\left(1-\omega^{\text{prim}}(p)p^{-s}\right)L(s,\omega^{\text{prim}}).$ Hence, from (4), (7), and Theorem 2.7, for any integer $m$ with $\frac{k+l-2}{2}<m<m+1<k$, it follows that:
@@ 2.2. An algorithm for the algebraic part of $D(m,f\times g)$
    Extend $f_{\rho}$ to a basis consisting of eigenforms $S_{k}(\Gamma_{1}(N))$ which we denote by $\mathcal{B}=\{f_{0}:=f_{\rho},f_{1},\dots,f_{n}\}$. (The dimension is $n+1$.)
@@
This is the desired value due to Lemma 2.4, Lemma 2.3, Theorem 2.5, and $\langle f_{0},f_{i}\rangle=0$ for $i=1,\dots,n$ and $\langle f^{\rho},f^{\rho}\rangle=\langle f,f\rangle$. The constant $c_{r}$ appears in Theorem 2.5.
@@
2.3. An algorithm for the algebraic part of $D(m,f)$
@@ 2.3.1. Special values of $L$-functions via modular symbols
$P,\gamma]:=\gamma(P\otimes\{0,\infty\})\in\mathbb{M}_{k}(\Gamma).$
When $\Gamma=\Gamma_{1}(N)$ and if $\gamma=\left(\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\right),\ \gamma^{\prime}=\left(\begin{smallmatrix}a^{\prime}&b^{\prime}\\ c^{\prime}&d^{\prime}\end{smallmatrix}\right)\in\Gamma_{1}(N)$ are such that $(c,d)\equiv(c^{\prime},d^{\prime})\pmod{N}$, then $P,\gamma]=[P,\gamma^{\prime}]$. Hence, the Manin symbol $P,\gamma]$ is determined by $P$ and the lower two entries $c,d$ of the matrix $\gamma$. So we take $[P,(c,d)]$ itself to be a Manin symbol. For the following proposition, see, for example, Stein [16, Prop. 8.3].
Proposition 2.8.
The Manin Symbols generate $\mathbb{M}_{k}(\Gamma).$
For any ring $R$ put $\mathbb{M}_{k}(\Gamma,R):=\mathbb{M}_{k}(\Gamma)\otimes_{\mathbb{Z}}R.$ There is a notion of boundary modular symbols and cuspidal symbols denoted by $\mathbb{B}_{k}(\Gamma)$ and $\mathbb{S}_{k}(\Gamma)$ respectively. One can also define Hecke operators on the space of modular symbols; see Stein [16, Section 8.3]. The space of cuspidal symbols is stable under the action of Hecke operators.
@@
Let $S_{k}(\Gamma)$ and $\bar{S}_{k}(\Gamma)$ denote the space of holomorphic and anti-holomorphic cusp forms respectively. Then the pairing $(\cdot,\cdot):(S_{k}(\Gamma)\oplus\bar{S}_{k}(\Gamma))\times\mathbb{M}_{k}(\Gamma,\mathbb{C})\rightarrow\mathbb{C}$
is Hecke equivariant, i.e., for $x\in\mathbb{M}_{k}(\Gamma,\mathbb{C})$ we have $(T_{n}(f_{1},f_{2}),x)=(f_{1},f_{2},T_{n}(x)).$
Assume $\Gamma$ is a finite index subgroup of $\mathrm{SL}_{2}(\mathbb{Z})$ such that $\eta\Gamma\eta=\Gamma,$ where $\eta=\text{diag}[-1,1].$ Then an involution $\iota^{*}$ on a Manin symbol is defined by $\iota^{*}[P(X,Y),(u,v)]=[P(X,-Y),(-u,v)].$ The involution $\iota^{*}$ commutes with the action of Hecke operators and it stabilizes the cuspidal subspace. We denote by $\mathbb{S}_{k}(\Gamma)^{-}$ and $\mathbb{S}_{k}(\Gamma)^{+}$ to be subspace of the cuspidal symbols which are the 1-1 and +1+1 eigenspaces of the action of $\iota^{*},$ respectively.
@@
The integration pairing $(\cdot,\cdot)$ induces nondegenerate Hecke-equivariant bilinear pairings
${S}_{k}(\Gamma)\times\mathbb{S}_{k}(\Gamma,\mathbb{C})^{-}\rightarrow\mathbb{C}\quad\text{and}\quad\overline{S}_{k}(\Gamma)\times\mathbb{S}_{k}(\Gamma,\mathbb{C})^{+}\rightarrow\mathbb{C}.$
The signs $\pm$ are interchanged in the above theorem in comparison with the one given in [16] because we do not take the ‘-’ sign in the definition of $\iota^{*}.$ This convention is adopted because it is the one implemented in SAGE [13] which we use for the computations that follow.
For a cusp form $f\in S_{k}(N)$ with Fourier expansion $\sum a(n,f)q^{n}$, the associated Dirichlet series $D(s,f)$ and completed $L$-function $L(s,f)$ are given by:
(9) $D(s,f):=\sum_{n=1}^{\infty}\frac{a(n,f)}{n^{s}},\quad L(s,f):=(2\pi)^{-s}\Gamma(s)D(s,f)=\int_{0}^{\infty}f(\sqrt{-1}t)t^{s}\frac{dt}{t},\quad\Re(s)\gg 0.$
Consider the modular symbols $X^{j}Y^{k-2-j}\otimes\{0,\infty\}$ for $j=0,1,2,\dots,k-2$. The non-degenerate and Hecke equivariant pairing of these modular symbols against the cusp form $f$ is:
@@
By (9), for $m=1,2\dots,k-1$, we have:
@@
For $m=1,2\dots,k-1$, define the algebraic part of a critical value $L(m,f)$ by:
@@
2.3.2. Some preliminaries for $\mathrm{SL}_{2}(\mathbb{Z})$
The following are well known:
  1. (1)
    The space of cusp forms $S_{k}(\mathrm{SL}_{2}(\mathbb{Z}))$ has co-dimension $1$ in $M_{k}(\mathrm{SL}_{2}(\mathbb{Z}))$ and let $d$ denote the dimension of $S_{k}(\mathrm{SL}_{2}(\mathbb{Z}))$.
  2. (2)
    The space $\mathbb{M}_{k}(\mathrm{SL}_{2}(\mathbb{Z}),\mathbb{Q})^{-}$ (resp., $\mathbb{M}_{k}(\mathrm{SL}_{2}(\mathbb{Z}),\mathbb{Q})^{+}$) is generated by $X^{i}Y^{k-2-i}\otimes\{0,\infty\}$, or equivalently by the Manin symbols $X^{i}Y^{k-2-i},(0,1)]$ for odd $i$ (resp., for even $i$). For brevity, we denote the Manin symbol $X^{i}Y^{k-2-i},(0,1)]$ by $(i,0,1)$.
  3. (3)
    $\mathbb{S}_{k}(\mathrm{SL}_{2}(\mathbb{Z}),\mathbb{Q})^{-}=\mathbb{M}_{k}(\mathrm{SL}_{2}(\mathbb{Z}),\mathbb{Q})^{-}.$
  4. (4)
    $\mathbb{S}_{k}(\mathrm{SL}_{2}(\mathbb{Z}),\mathbb{Q})^{+}$ has co-dimension $1$ in $\mathbb{M}_{k}(\mathrm{SL}_{2}(\mathbb{Z}),\mathbb{Q})^{+}.$
2.3.3.
The algorithm for even critical points.
Input: The first $d$ Fourier coefficients $a(1,f)=1,\,a(2,f),\dots,\,a(d,f)$ of a normalized eigenform $f\in S_{k}(\mathrm{SL}_{2}(\mathbb{Z}))$ which are assumed to be real.
Output: $\frac{D(m,f)}{(-2\pi\sqrt{-1})^{m-2}D(2,f)}$ for $m=2,4,\dots,k-2.$